# Tales of stellar and binary co-evolution, told by stellar oscillations and age estimates in main-sequence and red-giant stars

### v1

@@ Abstract
Aims. We aim to assess a holistic picture of the binary population and evolution in the red giant phase through characterizing the sample of binaries hosting oscillating red giants from the combination of large asteroseismic, spectroscopic, and astrometric surveys.
Methods. We investigate the binary properties of evolved stars in the APOKASC 3 and APO-K2 catalogs, leveraging asteroseismic constraints and Gaia DR3 non-single-star solutions. We explore the mass distribution of red-giant binary systems, analyze the evolution of their binary fraction. We investigate the impact of stellar evolution on the orbital periods (), eccentricities, radial velocity amplitudes and the fractional radius and identify candidate systems that may have undergone significant interactions.
Results. For stars with M 1.8M ⊙, we find binary fractions 31% and 41% for oscillating and non-oscillating solar-like stars on the main-sequence (MS). By using the peak frequency of the oscillation power excess () as luminosity proxy and evolutionary states, we detect a binary attrition of 69% and 81% on the low- and high-luminosity red-giant branch (RGB) and an additional 38% to the red clump (RC), with respect to the MS. Binaries hosting RC and secondary clump stars (2RC) stars are largely depleted at 500 and 200 days, respectively. We identify a population of rapidly rotating RC stars in short-period orbits as potential post-common-envelope merger products. Mass-dependent differences in binary fractions and orbital properties point to stronger binary attrition for stars with M 1.8 M⊙.
Conclusions. Binarity is not the primary cause of reduced oscillation amplitudes in main-sequence solar-like stars. The distinct mass distributions and depletion of short-period binaries in the red-giant phase underscore the impact of stellar expansion and binary interaction on stellar evolution. RC systems with 800 to 1,000 days are likely shaped by past interactions, such as mass transfer or loss, which can lead to significantly biased age estimates if not accounted for. This demonstrates the importance of identifying stellar binarity when using red giants as population tracers.
@@
Unless formed through a rare capturing event, binary stars are generally coeval, having formed from the same molecular cloud (Offner et al., 2023), justifying the assumption of an identical initial chemical composition (Torres et al., 2010). Therefore, the ages of its stellar components are typically assumed to be similar, within the free-fall time of the star-forming filament of the interstellar cloud from which the binary system has emerged (Offner et al., 2023). At a common distance, differences in luminosity and color can be attributed to differences in evolutionary stage rather than interstellar extinction. Given these shared characteristics, the key parameter governing the disparity between the two components is their mass difference, which can be accurately determined from the radial velocity (RV) solution of a double-lined spectroscopic binary (SB2) system. In particular, in the red-giant phase, mass ratios differing from unity by a few percent will lead to pronounced differences in the stellar evolution of the two components (Miglio et al., 2014; Beck et al., 2018a). Such systems are benchmarks for calibrating stellar models, testing complex physics, and validating model-dependent age estimates (del Burgo & Allende Prieto, 2018; Beck et al., 2018a; Jørgensen et al., 2020; Grossmann et al., 2025).
@@
Asteroseismology of solar-like oscillators has proven to be a powerful tool to characterize individual stars and stellar ensembles. Solar-like oscillations are stochastically excited by convective motions in the outer layers of stars and are observed from the main sequence through to the red-giant branch (e.g. Chaplin & Miglio, 2013; Houdek & Dupret, 2015, and references therein). Their frequencies are susceptible to the stellar structure, enabling the determination of key parameters such as mass, radius (Kjeldsen & Bedding, 1995; Kallinger et al., 2010), and age (e.g., Creevey et al., 2017; Anders et al., 2023). The discovery of mixed modes (Beck et al., 2011; Bedding et al., 2011) extended the seismic analysis into the central regions of the star, allowing for the determination of the evolutionary state (Bedding et al., 2011; Mosser et al., 2014) and constraining the internal rotational gradient (Beck et al., 2012). Binary systems play an essential role in seismology for calibrating the scaling relations (Gaulme et al., 2016; Benbakoura et al., 2021; Beck et al., 2025) and stellar models (Grossmann et al., 2025; Schimak et al., 2025; Thomsen et al., 2025).
@@
In this paper, we investigate the evolution of the binary fraction, orbital properties, and mass distribution of red-giant stars in binary systems by leveraging the combination of asteroseismic constraints and Gaia DR3 binary solutions. We analyze samples of oscillating red giants from the APOKASC 3 (Pinsonneault et al., 2025) and APO-K2 (Zinn et al., 2022) catalogs, supplemented with homogeneous spectroscopic parameters from APOGEE (Apache Point Observatory Galactic Evolution Experiment; Majewski et al., 2017) (Sect. 2). By cross-matching these samples with the NSS orbital solutions and binary indicators, we study how the binary fraction varies with stellar mass and evolutionary state (Sect. 3), and explore the impact of binarity on the red-giant mass distribution (Sect. 4). In Sect. 5, we examine the attenuation of radial velocity amplitudes, which may indicate common-envelope evolution or merger events. Our findings provide new empirical constraints on the co-evolution of stellar and orbital properties in binary systems hosting evolved stars. Conclusions are presented in Sect. 6.
@@ 2 Sample definition
The assumption that the photometrically dominating star is the primary does not hold for systems where the primary has already reached the He-core burning red-clump phase, while the secondary still progresses on the RGB, probably more luminous than the primary. In such systems, the mass ratio may significantly deviate from the initial value due to increased mass loss at the RGB tip, and it can even be inverted (Schimak et al., 2025). Given the brief duration of the red-giant phase, such systems are expected to be rare (Miglio et al., 2014; Mazzi et al., 2025).
@@ 2.1 APOKASC 3 catalog
APOKASC 3 presents the seismically inferred evolutionary state of the oscillating red giant based on the dipole-mixed-mode period-spacing measurements () presented by Vrard et al. (2025). This parameter distinguishes the evolutionary phase, specifically RGB and He-core burning stars (Bedding et al., 2011; Mosser et al., 2014). However, the catalog does not discriminate between red-clump (RC) stars and more massive secondary-clump (2RC) stars, which both are in the phase of quiescent He-core burning but differ in whether helium ignition occurred under degenerate or non-degenerate conditions. In our analysis, we distinguish them based on their seismically inferred masses, using 1.8 M⊙ as the threshold. The color code in Fig. 1 discriminates between these three evolutionary states in the giant phase.
The distinction between RGB and AGB stars becomes increasingly uncertain for more luminous stars than the RC. As noted by Pinsonneault et al. (2025), there is approximately a one-in-six chance that a star identified as being on the RGB is, in fact, on the AGB. Seismic discrimination between these two evolutionary phases remains challenging (Dréau et al., 2021). Following Pinsonneault et al. (2025), we treat AGB stars like RGB stars throughout our analysis. Using seismic parameters, inferred masses and evolutionary states were also used to obtain the stellar ages from comparison with single-star models.
@@ 2.2 APO-K2 catalog
The third version of the APO-K2 Catalog222The APO-K2 catalog files are electronically available for download from CDS through the reference J/ApJ/926/191. by Zinn et al. (2022) presents a similarly comprehensive study, using APOGEE spectroscopy and asteroseismic detections of 7,673 oscillating red-giant stars from the Campaigns C1-C8 and C10-18 of the K2 mission. As this work predates Gaia DR3, it uses DR2 parallaxes and a different correction method for . Furthermore, spectroscopic values from the APOGEE DR16 and GALAH DR2 are used. Due to the shorter time base and observational systematics of the K2 mission introduced by the satellite repositioning every 6 hours, stars in the APO-K2 sample are found in a smaller range of oscillation frequencies, 200 [Hz] 5 (Fig. 1). For this sample, the median fractional uncertainties in and are better than 2% and 1%, respectively. The fractional uncertainties for mass, radius, and age 11%, 7%, and 20%, respectively.
@@ 2.3 Binary sample selection
As source of binarity indicators, we use the non-single-star catalog (NSS, Gaia Collaboration et al., 2023a), published as part of the third data release (DR3) of the ESA Gaia mission (Gaia Collaboration et al., 2016, 2023b). It provides several sub-catalogs. The Gaia Collaboration cautions that the selection effects of the NSS catalog are difficult to quantify due to the heterogeneous detection algorithms used in its processing pipeline (Halbwachs et al., 2023). Recent studies have started to address this challenge by modeling the detection probabilities of astrometric binaries (El-Badry et al., 2024; Lam et al., 2025).
First, we use the catalog of two-body orbit solutions (TBO), which presents about half a million orbital parameters derived from the RV time series, obtained with the Gaia Radial Velocity Spectrometer (RVS, Katz et al., 2023), astrometry (Holl et al., 2023; Halbwachs et al., 2023) and photometry (Mowlavi et al., 2023). Gaia Collaboration et al. (2023a) point out that these are binary candidates for which completeness was chosen over purity. We follow a similar approach and do not filter solutions based on Gaia DR3 significance. From a detailed comparison of the SB9 catalog (Ninth Catalog of Spectroscopic Binary Orbits; Pourbaix et al., 2004), Beck et al. (2024) showed a good agreement for orbital periods Porb 1,500 days. We adopt a solution from the TBO as a binary candidate if at least the orbital period is provided in the catalog. For stars for which a single star model can not sufficiently explain the proper motion and no spectroscopic solution for an orbit was found, Gaia Collaboration et al. (2023a) provides non-linear or acceleration solutions (hereafter NLAC). The comparison with the SB9 confirmed that NLAC solutions indicate binaries with orbital periods typically much longer than 1,000 days (Beck et al., 2024).
@@
2.4 Rapid rotating red giants
Finally, we focus on the rapid rotators subsample in the APOKASC 3 sample. In a previous study, Patton et al. (2024) has identified 1500 active red giant candidates from the full APOGEE DR16, whose projected rotation velocity exceeds 5 km/s. For the stars with asteroseismic solutions from Kepler observations, Pinsonneault et al. (2025) determined mass, radius, and the evolutionary state for 332 objects. We include this subsample, depicted in Fig. 2, as enhanced rotation could be the signature of merger products or post-common envelope phases (Kochanek et al., 2014). Table 7 provides the parameters of the 14 systems that host rapidly rotating giants.
@@ 3 Evolution of binary fraction
Systems with orbital periods that are not sufficiently wide to allow the giant to expand freely until the tip of the RGB will undergo a common-envelope (CE) phase, likely leading to a short-lived destructive interaction. Consequently, the binary fraction is expected to decrease between the main sequence and the RGB and RC phases. However, the lack of consistent and complete binary indicators for large ensembles in the past made it difficult to estimate the evolutionary depletion of binary systems. From their large-sample study of RV variations in APOGEE data Badenes et al. (2018), they reported a 60% reduction from MS to RGB, with an estimated uncertainty of 20%.
The determination of the binary fraction relies on stellar multiplicity indicators rather than on the full set of orbital parameters for each system. Therefore, we use the BUF from Godoy-Rivera et al. (2025, see Table 1), which offers the most comprehensive compilation of all available binary indicators for stars in the Kepler field of view, from Gaia with RUWE (33.8%), RV Variable (5.7%), NSS (5.6%), Eclipsing Kepler (4.0%), Eclipsing Gaia (1.2%), as well as entries from ground-based surveys, like the SB9 (0.1%), binary stars from the NASA exoplanet archive (0.1%), indications from the HIPPARCOS-Gaia catalog of accelerations (HGCA) (0.1%), and entries from the Washington double star (WDS) catalog (4.0%). We excluded the Gaia Variablity flag (1.1%). An additional 44.2% of stars were found in more than one of these categories. As such, compilation of binary markers is not yet available for the APO-K2 sample, we only compute the binary fraction from the APOKASC 3 sample, as combining both samples would lead to a skewed result due to the differing completeness factors between the two catalogs.
The rich, homogeneous asteroseismic and spectroscopic dataset of the APOKASC 3 catalog provides an ideal opportunity to investigate the reduction in the binary fraction with advancing stellar evolution. We construct a magnitude-limited sample to determine a robust estimate of the binary fraction. The binary detection using Gaia shows an explicit cutoff with G 13 mag (Gaia Collaboration et al., 2023a), which is brighter than the faint and well-populated end of the sample of oscillating red giants from the Kepler mission (G 15 mag). We, therefore, adopt 13 mag as the faint and Gaia’s saturation limit of 4 mag as the bright cutoff of our sample. Table 5 reports the number of oscillating stars categorized by the seismically inferred evolutionary state of the asteroseismic primary.
@@ 3.1 Evolutionary subsample definition
Depending on its mass, a red-giant star can undergo two very distinct evolutionary channels (for details see the monographs by, e.g. Kippenhahn et al., 2013; Salaris & Cassisi, 2005; Pinsonneault & Ryden, 2023). The masses of giants in the APOKASC 3 catalog are dominated by low-mass stars but extend up to 3 M⊙ (Fig. 2). If the stellar mass is M/M⊙ 1.8, the inert helium core degenerates fully before reaching the ignition temperature of He. As a result, the star must ascend the RGB and continue to expand while the core heats up. With the onset of quiescent He-core burning at the tip of the RGB (up to R/R⊙ 200 1 AU), following a series of off-center helium flashes, the star settles into the RC phase (Bildsten et al., 2012). Stars more massive than this limit (M/M⊙ 1.8) ignite their helium core before it degenerates, doing so at significantly smaller radii on the RGB (200 R/R⊙ 80), and subsequently settle into the 2RC.
Furthermore, we must consider the differing binary fractions of progenitors of the red-giant stars in the sample. On the main sequence, the binary fraction is a strong function of the mass of the primary component (Lada, 2006). The main-sequence progenitors of these more massive stars are early to mid-A-type stars. While 40 - 50% of all solar-like dwarf stars (G and mid F) are found to be in binary systems, this fraction increases to 50 - 70% for the more massive early F and A-type stars on the main sequence (e.g., Offner et al., 2023; Moe & Di Stefano, 2017).
The precisely seismically inferred masses allow for an effective separation between the degenerate and non-degenerate evolutionary channels in the red-giant phase, enabling tests for these population differences. We, therefore, evaluate the binary fraction separately for these two evolutionary regimes (M 1.8 M⊙).

### v2

# Tales of stellar and binary coevolution, told by stellar oscillations and age estimates in main-sequence and red-giant stars

@@ Abstract
Aims. We aim to assess a holistic picture of the binary population and evolution in the red-giant phase by characterizing the sample of binaries hosting oscillating red giants from a combination of large asteroseismic, spectroscopic, and astrometric surveys.
Methods. We investigated the binary properties of evolved stars in the APOKASC 3 and APO-K2 catalogs, leveraging asteroseismic constraints and Gaia DR3 non-single-star solutions. We explored the mass distribution of red-giant binary systems, analyzed the evolution of their binary fraction. We investigated the impact of stellar evolution on the orbital periods (), eccentricities, radial velocity (RV) amplitudes, and the fractional radius and identifed candidate systems that may have undergone significant interactions.
Results. For stars with M 1.8M ⊙, we find binary fractions 31% and 41% for oscillating and non-oscillating solar-like stars on the main sequence (MS). By using the peak frequency of the oscillation power excess () as luminosity proxy and evolutionary states, we detect a binary attrition of 69% and 81% on the low- and high-luminosity red-giant branch (RGB) and an additional 38% to the red clump (RC), with respect to the MS. Binaries hosting RC and secondary-clump stars (2RC) are largely depleted at 500 and 200 days, respectively. We identify a population of rapidly rotating RC stars in short-period orbits as potential post-common-envelope merger products. Mass-dependent differences in binary fractions and orbital properties point to stronger binary attrition for stars with M 1.8 M⊙.
Conclusions. Binarity is not the primary cause of reduced oscillation amplitudes in MS solar-like stars. The distinct mass distributions and depletion of short-period binaries in the red-giant phase underscore the impact of stellar expansion and binary interaction on stellar evolution. Helium-core burning RC systems with 800 to 1,000 days are likely shaped by past interactions, such as mass transfer or loss, which can lead to significantly biased age estimates if not accounted for. This demonstrates the importance of identifying stellar binarity when using red giants as population tracers.
@@
Unless formed through a rare capturing event, binary stars are generally coeval, having formed from the same molecular cloud (Offner et al., 2023), justifying the assumption of an identical initial chemical composition (Torres et al., 2010). Therefore, the ages of its stellar components are typically assumed to be similar, within the free-fall time of the star-forming filament of the interstellar cloud from which the binary system has emerged (Offner et al., 2023). At a common distance, differences in luminosity and color can be attributed to differences in evolutionary stage rather than interstellar extinction. Given these shared characteristics, the key parameter governing the disparity between the two components is their mass difference, which can be accurately determined from the radial velocities (RV) solution of a double-lined spectroscopic binary (SB2) system. In particular, in the red-giant phase, mass ratios differing from unity by a few percent leads to pronounced differences in the stellar evolution of the two components (Miglio et al., 2014; Beck et al., 2018a). Such systems are benchmarks for calibrating stellar models, testing complex physics, and validating model-dependent age estimates (del Burgo & Allende Prieto, 2018; Beck et al., 2018a; Jørgensen et al., 2020; Grossmann et al., 2025).
@@
Asteroseismology of solar-like oscillators has proven to be a powerful tool to characterize individual stars and stellar ensembles. Solar-like oscillations are stochastically excited by convective motions in the outer layers of stars and are observed from the main sequence (MS) through to the red-giant branch (RGB) (e.g., Chaplin & Miglio, 2013; Houdek & Dupret, 2015, and references therein). Their frequencies are susceptible to the stellar structure, enabling the determination of key parameters such as mass, radius (Kjeldsen & Bedding, 1995; Kallinger et al., 2010), and age (e.g., Creevey et al., 2017; Anders et al., 2023). The discovery of mixed modes (Beck et al., 2011; Bedding et al., 2011) extended the seismic analysis into the central regions of the star, allowing for the determination of the evolutionary state (Bedding et al., 2011; Mosser et al., 2014) and constraining the internal rotational gradient (Beck et al., 2012). Binary systems play an essential role in seismology for calibrating the scaling relations (Gaulme et al., 2016; Benbakoura et al., 2021; Beck et al., 2026) and stellar models (Grossmann et al., 2025; Schimak et al., 2026; Thomsen et al., 2025).
@@
In this paper, we investigate the evolution of the binary fraction, orbital properties, and mass distribution of red-giant stars in binary systems by leveraging the combination of asteroseismic constraints and Gaia DR3 binary solutions. We analyze samples of oscillating red giants from the APOKASC 3 (Pinsonneault et al., 2025) and APO-K2 (Zinn et al., 2022) catalogs, supplemented with homogeneous spectroscopic parameters from APOGEE (Apache Point Observatory Galactic Evolution Experiment; Majewski et al., 2017) (Sect. 2). By cross-matching these samples with the NSS orbital solutions and binary indicators, we study how the binary fraction varies with stellar mass and evolutionary state (Sect. 3), and explore the impact of binarity on the red-giant mass distribution (Sect. 4). In Sect. 5, we examine the attenuation of RV amplitudes, which may indicate common-envelope evolution or merger events. Our findings provide new empirical constraints on the coevolution of stellar and orbital properties in binary systems hosting evolved stars. Conclusions are presented in Sect. 6.
@@ 2 Sample definition
The assumption that the photometrically dominating star is the primary does not hold for systems where the primary has already reached the He-core burning red-clump (RC) phase, while the secondary still progresses on the RGB, probably more luminous than the primary. In such systems, the mass ratio may significantly deviate from the initial value due to increased mass loss at the RGB tip, and it can even be inverted (Schimak et al., 2026). Given the brief duration of the red-giant phase, such systems are expected to be rare (Miglio et al., 2014; Mazzi et al., 2025).
@@ 2.1 APOKASC 3 catalog
APOKASC 3 presents the seismically inferred evolutionary state of the oscillating red giant based on the dipole-mixed-mode period-spacing measurements () presented by Vrard et al. (2025). This parameter distinguishes the evolutionary phase, specifically RGB and He-core burning stars (Bedding et al., 2011; Mosser et al., 2014). However, the catalog does not discriminate between RC stars and more massive secondary-clump (2RC) stars, which both are in the phase of quiescent He-core burning but differ in whether helium ignition occurred under degenerate or non-degenerate conditions. In our analysis, we distinguish them based on their seismically inferred masses, using 1.8 M⊙ as the threshold. The color code in Fig. 1 discriminates between these three evolutionary states in the giant phase.
The distinction between RGB and AGB stars becomes increasingly uncertain for more luminous stars than the RC. As is noted by Pinsonneault et al. (2025), there is approximately a one-in-six chance that a star identified as being on the RGB is, in fact, on the AGB. Seismic discrimination between these two evolutionary phases remains challenging (Dréau et al., 2021). Following Pinsonneault et al. (2025), we treat AGB stars like RGB stars throughout our analysis. Using seismic parameters, inferred masses and evolutionary states were also used to obtain the stellar ages from comparison with single-star models.
@@ 2.2 APO-K2 catalog
The third version of the APO-K2 Catalog222The APO-K2 catalog files are electronically available for download from CDS through the reference J/ApJ/926/191. by Zinn et al. (2022) presents a similarly comprehensive study, using APOGEE spectroscopy and asteroseismic detections of 7,673 oscillating red-giant stars from the Campaigns C1-C8 and C10-18 of the K2 mission. As this work predates Gaia DR3, it uses DR2 parallaxes and a different correction method for . Furthermore, spectroscopic values from the APOGEE DR16 and GALAH DR2 are used. Due to the shorter time base and observational systematics of the K2 mission introduced by the satellite repositioning every 6 hours, stars in the APO-K2 sample are found in a smaller range of oscillation frequencies, 200 [Hz] 5 (Fig. 1). For this sample, the median fractional uncertainties in and are better than 2% and 1%, respectively. The fractional uncertainties for mass, radius, and age are 11%, 7%, and 20%, respectively.
@@ 2.3 Binary sample selection
As a source of binarity indicators, we used the non-single-star catalog (NSS, Gaia Collaboration et al., 2023a), published as part of the third data release (DR3) of the ESA Gaia mission (Gaia Collaboration et al., 2016, 2023b). It provides several sub-catalogs. The Gaia Collaboration cautions that the selection effects of the NSS catalog are difficult to quantify due to the heterogeneous detection algorithms used in its processing pipeline (Halbwachs et al., 2023). Recent studies have started to address this challenge by modeling the detection probabilities of astrometric binaries (El-Badry et al., 2024; Lam et al., 2025).
First, we used the catalog of two-body orbit solutions (TBO), which presents about half a million orbital parameters derived from the RV time series, obtained with the Gaia Radial Velocity Spectrometer (RVS, Katz et al., 2023), astrometry (Holl et al., 2023; Halbwachs et al., 2023) and photometry (Mowlavi et al., 2023). Gaia Collaboration et al. (2023a) point out that these are binary candidates for which completeness was chosen over purity. We followed a similar approach and did not filter solutions based on Gaia DR3 significance. From a detailed comparison of the SB9 catalog (Ninth Catalog of Spectroscopic Binary Orbits; Pourbaix et al., 2004), Beck et al. (2024) showed a good agreement for orbital periods Porb 1,500 days. We adopt a solution from the TBO as a binary candidate if at least the orbital period is provided in the catalog. For stars for which a single star model cannot sufficiently explain the proper motion and no spectroscopic solution for an orbit was found, Gaia Collaboration et al. (2023a) provides nonlinear or acceleration solutions (hereafter NLAC). The comparison with the SB9 confirmed that NLAC solutions indicate binaries with orbital periods typically much longer than 1,000 days (Beck et al., 2024).
@@
2.4 Rapidly rotating red giants
Finally, we focus on the rapid rotators subsample in the APOKASC 3 sample. In a previous study, Patton et al. (2024) has identified 1500 active red-giant candidates from the full APOGEE DR16, whose projected rotation velocity exceeds 5 km/s. For the stars with asteroseismic solutions from Kepler observations, Pinsonneault et al. (2025) determined mass, radius, and the evolutionary state for 332 objects. We include this subsample, depicted in Fig. 2, as enhanced rotation could be the signature of merger products or post-common envelope phases (Kochanek et al., 2014). Table 7 provides the parameters of the 14 systems that host rapidly rotating giants.
@@ 3 Evolution of binary fraction
Systems with orbital periods that are not sufficiently wide to allow the giant to expand freely until the tip of the RGB will undergo a common-envelope (CE) phase, likely leading to a short-lived destructive interaction. Consequently, the binary fraction is expected to decrease between the MS and the RGB and RC phases. However, the lack of consistent and complete binary indicators for large ensembles in the past made it difficult to estimate the evolutionary depletion of binary systems. From their large-sample study of RV variations in APOGEE data Badenes et al. (2018), they reported a 60% reduction from MS to RGB, with an estimated uncertainty of 20%.
The determination of the binary fraction relies on stellar multiplicity indicators rather than on the full set of orbital parameters for each system. Therefore, we used the BUF from Godoy-Rivera et al. (2025, see Table 1), which offers the most comprehensive compilation of all available binary indicators for stars in the Kepler field of view, from Gaia with RUWE (33.8%), RV Variable (5.7%), NSS (5.6%), Eclipsing Kepler (4.0%), Eclipsing Gaia (1.2%), as well as entries from ground-based surveys, like the SB9 (0.1%), binary stars from the NASA exoplanet archive (0.1%), indications from the HIPPARCOS-Gaia catalog of accelerations (HGCA) (0.1%), and entries from the Washington double star (WDS) catalog (4.0%). We excluded the Gaia Variablity flag (1.1%). An additional 44.2% of stars were found in more than one of these categories. As such, compilation of binary markers is not yet available for the APO-K2 sample, we only compute the binary fraction from the APOKASC 3 sample, as combining both samples would lead to a skewed result due to the differing completeness factors between the two catalogs.
The rich, homogeneous asteroseismic and spectroscopic dataset of the APOKASC 3 catalog provides an ideal opportunity to investigate the reduction in the binary fraction with advancing stellar evolution. We constructed a magnitude-limited sample to determine a robust estimate of the binary fraction. The binary detection using Gaia shows an explicit cutoff with G 13 mag (Gaia Collaboration et al., 2023a), which is brighter than the faint and well-populated end of the sample of oscillating red giants from the Kepler mission (G 15 mag). We therefore adopted 13 mag as the faint and Gaia’s saturation limit of 4 mag as the bright cutoff of our sample. Table 5 reports the number of oscillating stars categorized by the seismically inferred evolutionary state of the asteroseismic primary.
@@ 3.1 Evolutionary subsample definition
Depending on its mass, a red-giant star can undergo two very distinct evolutionary channels (for details see the monographs by, e.g., Kippenhahn et al., 2013; Salaris & Cassisi, 2005; Pinsonneault & Ryden, 2023). The masses of giants in the APOKASC 3 catalog are dominated by low-mass stars but extend up to 3 M⊙ (Fig. 2). If the stellar mass is M/M⊙ 1.8, the inert helium core degenerates fully before reaching the ignition temperature of He. As a result, the star must ascend the RGB and continue to expand while the core heats up. With the onset of quiescent He-core burning at the tip of the RGB (up to R/R⊙ 200 1 AU), following a series of off-center helium flashes, the star settles into the RC phase (Bildsten et al., 2012). Stars more massive than this limit (M/M⊙ 1.8) ignite their helium core before it degenerates, doing so at significantly smaller radii on the RGB (200 R/R⊙ 80), and subsequently settle into the 2RC.
Furthermore, we must consider the differing binary fractions of progenitors of the red-giant stars in the sample. On the MS, the binary fraction is a strong function of the mass of the primary component (Lada, 2006). The main-sequence progenitors of these more massive stars are early to mid-A-type stars. While 40 - 50% of all solar-like dwarf stars (G and mid F) are found to be in binary systems, this fraction increases to 50 - 70% for the more massive early F- and A-type stars on the MS (e.g., Offner et al., 2023; Moe & Di Stefano, 2017).
The precisely seismically inferred masses allow for an effective separation between the degenerate and non-degenerate evolutionary channels in the red-giant phase, enabling tests for these population differences. We therefore evaluated the binary fraction separately for these two evolutionary regimes (M 1.8 M⊙).